OpenStudy (anonymous):

Help Please. I need it to be explained Solve the equation if possible, then write the irrational numbers in simplest radical form. x²-6x+1=0

OpenStudy (anonymous):

Use quadratic formula... x = [-b ± √b² - 4ac]/2a ax² + bx + c = 0 a = 1 b = -6 c = 1 x = [6 ± √(-6)² - 4(1)(1)]/2(1) x = [6 ± √36 - 4]/2 x = [6 ± √32]/2 x = [6 ± 4√2]/2 You get: x = 3 ± 2√2 I hope that helps!

OpenStudy (anonymous):

Thank you